Find the slope-intercept form of the line that passes through the given point and has the given slope. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.
(7, 1)
m = 10

Found 2 solutions by fractalier, Alan3354:
Answer by fractalier(6550)

Let us substitute your (x,y) and m into
y = mx + b
to find b...
1 = 10(7) + b
1 = 70 + b
b = -69
Now plug m and b back in and get
y = 10x - 69

Answer by Alan3354(69443)

Find the slope-intercept form of the line that passes through the given point and has the given slope. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.
(7, 1)
m = 10
------------
y-y1 = m*(x-x1)
y-1 = 10*(x-7)
===================
Solve for y to get slope-intercept form.